# Stochastic Stability of ACIMs for Piecewise Expanding $C^{1+\varepsilon}$ Maps

### 04 Apr 2026 v1

@@ 3. Perturbation Operators and Uniform Lasota–Yorke Inequality
Work on $\mathbb{T}$. Let $C_{1}:=2^{1/p}\max\bigl(1,\;4\|q^{\prime}\|_{\infty}\bigr)$. By Theorem 3 of [4], $\tau$ satisfies
$\mathrm{var}_{1,1/p}(P_{\tau}f)\;\leq\;\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}$
with $0<\alpha_{0}=s^{-1/p}+2s^{-1}<1/C_{1}$. Then there exist $\alpha\in(0,1)$ and $C>0$, independent of $\delta>0$, such that for all $f\in BV_{1,1/p}$,
@@
Let $g:=P_{\tau}f$, so $P_{\delta}f=Q_{\delta}g$ and $\|g\|_{1}=\|f\|_{1}$. By hypothesis,
(3.2) $\mathrm{var}_{1,1/p}(g)\;\leq\;\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}.$
We bound $\mathrm{var}_{1,1/p}(Q_{\delta}g)=\sup_{r>0}r^{-1/p}\mathrm{Osc}_{1}(Q_{\delta}g,r)$ by splitting at $r=\delta$.
Large scales ($r\geq\delta$). By Lemma 3.4, $\mathrm{Osc}_{1}(Q_{\delta}g,r)\leq\mathrm{Osc}_{1}(g,r+\delta)\leq(r+\delta)^{1/p}\mathrm{var}_{1,1/p}(g)$. Since $r+\delta\leq 2r$ when $r\geq\delta$,
$r^{-1/p}\mathrm{Osc}_{1}(Q_{\delta}g,r)\;\leq\;2^{1/p}\,\mathrm{var}_{1,1/p}(g).$
Small scales ($0<r<\delta$). By Lemma 3.3,
$\mathrm{Osc}_{1}(Q_{\delta}g,r)\;\leq\;\frac{Cr}{\delta}\,\mathrm{Osc}_{1}(g,\,r+\delta)\;\leq\;\frac{Cr}{\delta}\,(r+\delta)^{1/p}\,\mathrm{var}_{1,1/p}(g).$
Since $r<\delta$, the function $r\mapsto r^{1-1/p}(r+\delta)^{1/p}\delta^{-1}$ is nondecreasing (both factors increase in $r$ for $p\geq 1$), so its supremum on $(0,\delta)$ is attained at $r\nearrow\delta$:
$\sup_{0<r<\delta}\;r^{-1/p}\mathrm{Osc}_{1}(Q_{\delta}g,r)\;\leq\;C\cdot\frac{\delta^{1-1/p}(2\delta)^{1/p}}{\delta}\,\mathrm{var}_{1,1/p}(g)\;=\;2^{1/p}C\,\mathrm{var}_{1,1/p}(g).$
Combining.
$\mathrm{var}_{1,1/p}(Q_{\delta}g)\;\leq\;C_{1}\,\mathrm{var}_{1,1/p}(g),\qquad C_{1}:=\max\!\bigl(2^{1/p},\;2^{1/p}C\bigr)=2^{1/p}\max\!\bigl(1,\;4\|q^{\prime}\|_{\infty}\bigr).$
Substituting (3.2) and using $\|P_{\delta}f\|_{1}\leq\|f\|_{1}$,
$\|P_{\delta}f\|_{1,1/p}\;\leq\;\underbrace{C_{1}\alpha_{0}}_{=:\,\alpha}\|f\|_{1,1/p}+\underbrace{(C_{1}C_{0}+1)}_{=:\,C}\|f\|_{1}.$
Since $\alpha_{0}<1/C_{1}$, we have $\alpha\in(0,1)$, and both constants are independent of $\delta$. ∎
@@ 4. Existence and Convergence of Invariant Densities
In this section, we prove the existence of invariant densities for the perturbed operators $P_{\delta}$ and establish their convergence in $L^{1}$ to the invariant density of the unperturbed system.
@@
For each $\delta>0$, the operator $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}$ such that
@@
Let $f\in BV_{1,1/p}$ with $f\geq 0$ and $\|f\|_{1}=1$. By Theorem 3.5, there exist constants $\lambda\in(0,1)$ and $C>0$, independent of $\delta$, such that
$\|P_{\delta}g\|_{1,1/p}\leq\lambda\|g\|_{1,1/p}+C\|g\|_{1}\qquad\text{for all }g\in BV_{1,1/p}.$
Iterating and using $\|P_{\delta}^{k}f\|_{1}=\|f\|_{1}=1$ (since $f\geq 0$ and $P_{\delta}$ preserves the $L^{1}$ norm of nonnegative functions), we obtain for all $n\in\mathbb{N}$:
$\|P_{\delta}^{n}f\|_{1,1/p}\;\leq\;\lambda^{n}\|f\|_{1,1/p}+C\sum_{k=0}^{n-1}\lambda^{k}\|f\|_{1}\;\leq\;\|f\|_{1,1/p}+\frac{C}{1-\lambda}\;=:\;C^{\prime}.$
Define the Cesàro averages $h_{n}:=\frac{1}{n}\sum_{k=0}^{n-1}P_{\delta}^{k}f$. By convexity of the norm,
$\|h_{n}\|_{1,1/p}\;\leq\;\frac{1}{n}\sum_{k=0}^{n-1}\|P_{\delta}^{k}f\|_{1,1/p}\;\leq\;C^{\prime},$
so $\{h_{n}\}$ is bounded in $BV_{1,1/p}$. Since the embedding $BV_{1,1/p}\hookrightarrow L^{1}(I)$ is compact, there exist a subsequence $\{h_{n_{j}}\}$ and $h_{\delta}\in L^{1}(I)$ such that $h_{n_{j}}\to h_{\delta}$ in $L^{1}(I)$.
By lower semicontinuity of $\mathrm{var}_{1,1/p}$ with respect to $L^{1}$ convergence ([5], Proposition 2.3),
$\mathrm{var}_{1,1/p}(h_{\delta})\;\leq\;\liminf_{j\to\infty}\mathrm{var}_{1,1/p}(h_{n_{j}})\;\leq\;C^{\prime},$
so $h_{\delta}\in BV_{1,1/p}$.
@@
Taking $L^{1}$ norms and using $\|P_{\delta}^{n}f\|_{1}=\|f\|_{1}=1$:
@@
Since $P_{\delta}:L^{1}(I)\to L^{1}(I)$ is a bounded linear operator, $h_{n_{j}}\to h_{\delta}$ in $L^{1}$ implies $P_{\delta}h_{n_{j}}\to P_{\delta}h_{\delta}$ in $L^{1}$. Combined with $\|P_{\delta}h_{n_{j}}-h_{n_{j}}\|_{1}\to 0$ and $h_{n_{j}}\to h_{\delta}$, we conclude $P_{\delta}h_{\delta}=h_{\delta}$.
Finally, since $h_{n_{j}}\to h_{\delta}$ in $L^{1}$, there is a further subsequence along which $h_{n_{j}}(x)\to h_{\delta}(x)$ almost everywhere. Since each $h_{n_{j}}\geq 0$, we get $h_{\delta}\geq 0$ a.e. Moreover,
$\|h_{\delta}\|_{1}=\lim_{j\to\infty}\|h_{n_{j}}\|_{1}=1.\qed$
@@
$\sup_{g\in K}\|Q_{\delta}g-g\|_{1}\to 0\quad\text{as }\delta\to 0.$
@@
Since $K$ is relatively compact in $L^{1}(I)$, it is totally bounded. Hence, for any $\varepsilon>0$, there exist functions $g_{1},\dots,g_{N}\in L^{1}(I)$ such that for every $g\in K$ there exists $i$ with
@@
$\|Q_{\delta}g-g\|_{1}\leq 2\varepsilon+\|Q_{\delta}g_{i}-g_{i}\|_{1}.$
Taking the supremum over $g\in K$, we obtain
@@
For each fixed $i$, $\|Q_{\delta}g_{i}-g_{i}\|_{1}\to 0$ as $\delta\to 0$ (approximation of identity). Hence,
$\max_{1\leq i\leq N}\|Q_{\delta}g_{i}-g_{i}\|_{1}\to 0.$
@@
Assume $\tau$ admits a unique ACIM with density $h\in BV_{1,1/p}$. Let $h_{\delta}$ be an invariant density of $P_{\delta}$. Then $\|h_{\delta}-h\|_{1}\to 0$ as $\delta\to 0$.
@@
We define an equivalent norm on $BV_{1,1/p}$ by
$\|f\|_{\mathcal{B}}:=\mathrm{var}_{1,1/p}(f)+A\|f\|_{1},\qquad A>0\text{ chosen below},$
as the strong norm, and retain $\|\cdot\|_{L^{1}}:=\|\cdot\|_{1}$ as the weak norm.
@@
$\mathrm{var}_{1,1/p}(P_{\delta}f)\leq\alpha\,\mathrm{var}_{1,1/p}(f)+C\|f\|_{1},\qquad\|P_{\delta}f\|_{1}\leq\|f\|_{1}.$
Choose $A>C/(1-\alpha)$. Adding AA times the second inequality to the first:
$\|P_{\delta}f\|_{\mathcal{B}}\;\leq\;\alpha\|f\|_{\mathcal{B}}+B\|f\|_{L^{1}},\qquad B:=C+A(1-\alpha),$
uniformly in $\delta>0$. Since $\|f\|_{L^{1}}\leq A^{-1}\|f\|_{\mathcal{B}}$:
$\|P_{\delta}f\|_{\mathcal{B}}\;\leq\;M\|f\|_{\mathcal{B}},\qquad M:=\alpha+\tfrac{B}{A}=1+\tfrac{C}{A}<\infty,$
@@
Step 2: Perturbation smallness $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}}\to 0$. Write $P_{\delta}f-P_{\tau}f=(Q_{\delta}-I)P_{\tau}f$. Since $P_{\tau}$ satisfies the same Lasota–Yorke inequality [4], it maps the unit ball of $(BV_{1,1/p},\|\cdot\|_{\mathcal{B}})$ into a set bounded in $BV_{1,1/p}$. By compactness of $BV_{1,1/p}\hookrightarrow L^{1}(I)$, the set
@@
and hence $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}}\to 0$.
@@
Quasi-compactness. By the Lasota–Yorke inequality for $P_{\tau}$ and the compact embedding $BV_{1,1/p}\hookrightarrow L^{1}(I)$, the Ionescu–Tulcea–Marinescu theorem [8] implies that $P_{\tau}$ is quasi-compact on $(BV_{1,1/p},\|\cdot\|_{\mathcal{B}})$ with $r_{\mathrm{ess}}(P_{\tau})\leq\alpha<1$.
Simplicity of eigenvalue 11. Let $g\in BV_{1,1/p}$ with $P_{\tau}g=g$. Write $g=g^{+}-g^{-}$ with $g^{\pm}\geq 0$. By positivity of $P_{\tau}$:
$P_{\tau}|g|\;\geq\;|P_{\tau}g|\;=\;|g|.$
Since $P_{\tau}$ is a Markov operator, $\int P_{\tau}|g|\,dm=\int|g|\,dm$. Since $P_{\tau}|g|\geq|g|$ almost everywhere and both functions have the same integral, it follows that
$P_{\tau}|g|\;=\;|g|\quad\text{almost everywhere.}$
Adding this to $P_{\tau}g=g$:
@@
so $P_{\tau}g^{+}=g^{+}$ a.e., and likewise $P_{\tau}g^{-}=g^{-}$ a.e. If $g^{+}\not\equiv 0$, then $g^{+}/\|g^{+}\|_{1}$ is a normalized nonnegative fixed point of $P_{\tau}$; by uniqueness of the ACIM (see [4]), $g^{+}=c^{+}h$. Similarly $g^{-}=c^{-}h$. Hence $g=(c^{+}-c^{-})h$, and $\ker(P_{\tau}-I)$ is one-dimensional.
Isolation. Since $r_{\mathrm{ess}}(P_{\tau})\leq\alpha<1$, all eigenvalues of $P_{\tau}$ on the unit circle are isolated with finite multiplicity. The eigenvalue 11 is simple by the above, so it is isolated. Its spectral projection is $\Pi f=\bigl(\int f\,dm\bigr)h$, which has rank one.
Step 4: Application of Keller–Liverani. Steps 1–3 verify the hypotheses of [6, Theorem 1]:
@@
    Uniform Lasota–Yorke: $\|P_{\delta}f\|_{\mathcal{B}}\leq\alpha\|f\|_{\mathcal{B}}+B\|f\|_{L^{1}}$, with $\alpha\in(0,1)$ and $B<\infty$ independent of $\delta$.
@@
    Perturbation smallness: $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}}\to 0$.
@@
    Spectral isolation: eigenvalue 11 of $P_{\tau}$ is isolated and simple, with rank-one spectral projection $\Pi$.

### 11 Apr 2026 v2

@@ 3. Perturbation Operators and Uniform Lasota–Yorke Inequality
Remark 1.
When it is necessary to distinguish integration over the interval $I=[0,1]$ from integration over the torus $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, we write $\mathrm{Osc}_{1}^{I}(f,r)$ in place of $\mathrm{Osc}_{1}(f,r)$; the superscript indicates that the $L^{1}$-integration in (2.2) is taken over $I$. The unsuperscripted notation $\mathrm{Osc}_{1}(f,r)$ is used when the ambient space is clear from context. Explicit superscripts $I$ and $\mathbb{T}$ are introduced when both domains are considered simultaneously. In particular, equation (1.1) uses $\mathrm{Osc}_{1}^{I}$ to emphasize integration over the interval.
Lemma 3.5 (Interval-to-torus variation comparison).
For any $g\in BV_{1,1/p}(I)$,
$\operatorname{var}_{1,1/p}^{\,\mathbb{T}}(g^{\mathbb{T}})\;\leq\;(1+2^{1+1/p})\,\operatorname{var}_{1,1/p}^{\,I}(g),$
where $g^{\mathbb{T}}$ denotes the $\mathbb{Z}$-periodization of $g|_{I}$.
Proof.
It suffices to consider $r\in(0,1/2]$. For $r\geq 1/2$, the torus ball satisfies $B^{\mathbb{T}}(x,r)=\mathbb{T}$ for every $x\in\mathbb{T}$, so $\operatorname{Osc}_{1}^{\mathbb{T}}(g^{\mathbb{T}},r)$ is constant on $[1/2,1)$. Since $r^{-1/p}$ is strictly decreasing in $r$, the supremum defining $\operatorname{var}_{1,1/p}^{\mathbb{T}}(g^{\mathbb{T}})$ is therefore achieved at some $r\in(0,1/2]$, and it suffices to bound $r^{-1/p}\operatorname{Osc}_{1}^{\mathbb{T}}(g^{\mathbb{T}},r)$ for $r\leq 1/2$. For $r\leq 1/2$, split the torus oscillation integral:
(3.1) $\operatorname{Osc}_{1}^{\,\mathbb{T}}(g^{\mathbb{T}},r)=\int_{[r,1-r]}\operatorname{Osc}^{I}(g,r,x)\,dx+\int_{[0,r)\cup(1-r,1]}\operatorname{Osc}^{\mathbb{T}}(g^{\mathbb{T}},r,x)\,dx.$
The first integral is bounded by $\operatorname{Osc}_{1}^{I}(g,r)$.
For the boundary strips, if $x\in[0,r)$ the torus ball satisfies
$B^{\mathbb{T}}(x,r)=[0,x+r]\cup[1-r+x,1]\subset B^{I}(0,2r)\cup B^{I}(1,2r).$
Since $B^{\mathbb{T}}(x,r)\subset B^{I}(0,2r)\cup B^{I}(1,2r)$, oscillation on the boundary strip is controlled by the oscillations near the two endpoints, hence $\operatorname{Osc}^{\mathbb{T}}(g^{\mathbb{T}},r,x)\leq\operatorname{Osc}^{I}(g,2r,0)+\operatorname{Osc}^{I}(g,2r,1)$. The same bound holds for $x\in(1-r,1]$. Integrating over the boundary strips:
$\int_{[0,r)\cup(1-r,1]}\operatorname{Osc}^{\mathbb{T}}(g^{\mathbb{T}},r,x)\,dx\leq 2r\bigl[\operatorname{Osc}^{I}(g,2r,0)+\operatorname{Osc}^{I}(g,2r,1)\bigr].$
For $x_{0}\in\{0,1\}$, if $x_{1},x_{2}\in B^{I}(x_{0},2r)$ and $x\in B^{I}(x_{0},2r)$, then $|x-x_{1}|,|x-x_{2}|\leq 2r$, hence $|g(x_{1})-g(x_{2})|\leq\operatorname{Osc}^{I}(g,2r,x)$ a.e. Integrating over $x\in B^{I}(x_{0},2r)$:
$2r\cdot\operatorname{Osc}^{I}(g,2r,x_{0})\leq\int_{B^{I}(x_{0},2r)}\operatorname{Osc}^{I}(g,2r,x)\,dx\leq\operatorname{Osc}_{1}^{I}(g,2r).$
Therefore,
$\int_{[0,r)\cup(1-r,1]}\operatorname{Osc}^{\mathbb{T}}(g^{\mathbb{T}},r,x)\,dx\leq 2\,\operatorname{Osc}_{1}^{I}(g,2r).$
Substituting into (3.1) and multiplying by $r^{-1/p}$:
$r^{-1/p}\operatorname{Osc}_{1}^{\,\mathbb{T}}(g^{\mathbb{T}},r)$ $\leq r^{-1/p}\operatorname{Osc}_{1}^{I}(g,r)+2r^{-1/p}\operatorname{Osc}_{1}^{I}(g,2r)$
$\leq\operatorname{var}_{1,1/p}^{I}(g)+2\cdot 2^{1/p}(2r)^{-1/p}\operatorname{Osc}_{1}^{I}(g,2r)$
$\leq\operatorname{var}_{1,1/p}^{I}(g)+2^{1+1/p}\operatorname{var}_{1,1/p}^{I}(g)$
$=(1+2^{1+1/p})\operatorname{var}_{1,1/p}^{I}(g).$
Taking the supremum over $r\in(0,1/2]$ gives the result. ∎
@@
There exist constants $0<\alpha<1$ and $C>0$, independent of $\delta>0$, such that for all $f\in BV_{1,1/p}(I)$,
@@
Let $g:=P_{\tau}f$, so $P_{\delta}f=Q_{\delta}g$ and $\|g\|_{1}=\|f\|_{1}$. By Assumption 2.5:
(3.2) $\operatorname{var}^{I}_{1,1/p}(g)\leq\alpha_{0}\|f\|_{1,1/p}+C_{0}\|f\|_{1}.$
We estimate $\operatorname{var}_{1,1/p}^{I}(Q_{\delta}g)=\sup_{0<r\leq 1}r^{-1/p}\operatorname{Osc}_{1}^{I}(Q_{\delta}g,r)$.
Since $Q_{\delta}$ on $I$ equals the torus convolution $Q_{\delta}^{\mathbb{T}}$ restricted to $I$, we have $(Q_{\delta}g)(x)=(Q_{\delta}^{\mathbb{T}}g^{\mathbb{T}})(x)$ for all $x\in I$, and therefore
(3.3) $\operatorname{Osc}_{1}^{I}(Q_{\delta}g,r)\;\leq\;\operatorname{Osc}_{1}^{\,\mathbb{T}}(Q_{\delta}^{\mathbb{T}}g^{\mathbb{T}},r).$
We split into two cases.
Case 1: $r\geq\delta$. By Lemma 3.4 applied on $\mathbb{T}$:
$\operatorname{Osc}_{1}^{\,\mathbb{T}}(Q_{\delta}^{\mathbb{T}}g^{\mathbb{T}},r)\leq\operatorname{Osc}_{1}^{\,\mathbb{T}}(g^{\mathbb{T}},r+\delta)\leq(r+\delta)^{1/p}\operatorname{var}_{1,1/p}^{\,\mathbb{T}}(g^{\mathbb{T}}).$
Since $r+\delta\leq 2r$,
$r^{-1/p}\operatorname{Osc}_{1}^{I}(Q_{\delta}g,r)\;\leq\;2^{1/p}\operatorname{var}_{1,1/p}^{\,\mathbb{T}}(g^{\mathbb{T}}).$
Case 2: $0<r<\delta$. By Lemma 3.3 applied on $\mathbb{T}$:
$\operatorname{Osc}_{1}^{\,\mathbb{T}}(Q_{\delta}^{\mathbb{T}}g^{\mathbb{T}},r)\leq\frac{Cr}{\delta}(r+\delta)^{1/p}\operatorname{var}_{1,1/p}^{\,\mathbb{T}}(g^{\mathbb{T}}).$
Thus
$r^{-1/p}\operatorname{Osc}_{1}^{I}(Q_{\delta}g,r)\leq C\,r^{1-1/p}\delta^{-1}(r+\delta)^{1/p}\operatorname{var}_{1,1/p}^{\,\mathbb{T}}(g^{\mathbb{T}}).$
The function $\varphi(r)=r^{1-1/p}(r+\delta)^{1/p}$ is increasing on $(0,\delta)$, so $\sup_{0<r<\delta}\varphi(r)=\varphi(\delta)=2^{1/p}\delta$. Hence
$r^{-1/p}\operatorname{Osc}_{1}^{I}(Q_{\delta}g,r)\;\leq\;C\cdot 2^{1/p}\operatorname{var}_{1,1/p}^{\,\mathbb{T}}(g^{\mathbb{T}}).$
Combining both cases:
$\operatorname{var}_{1,1/p}^{I}(Q_{\delta}g)\;\leq\;C_{1}\operatorname{var}_{1,1/p}^{\,\mathbb{T}}(g^{\mathbb{T}}),$
where $C_{1}=2^{1/p}\max\{1,4\|q^{\prime}\|_{\infty}\}$, coming from Lemmas 3.3 and  3.4. Now apply Lemma 3.5:
$\operatorname{var}_{1,1/p}^{I}(Q_{\delta}g)\;\leq\;C_{1}(1+2^{1+1/p})\operatorname{var}_{1,1/p}^{I}(g).$
Recalling $\widetilde{C}_{1}=C_{1}(1+2^{1+1/p})$ from Assumption 3.1. By (3.2) and the $L^{1}$-contraction $\|P_{\delta}f\|_{1}\leq\|f\|_{1}$:
$\|P_{\delta}f\|_{1,1/p}\;\leq\;\widetilde{C}_{1}\alpha_{0}\|f\|_{1,1/p}+(\widetilde{C}_{1}C_{0}+1)\|f\|_{1}.$
Let $\alpha:=\widetilde{C}_{1}\alpha_{0}$ and $C:=\widetilde{C}_{1}C_{0}+1$. Both constants are independent of $\delta$. Since $\alpha_{0}<1/\widetilde{C}_{1}$ (Assumption 3.1), it follows that $\alpha=\widetilde{C}_{1}\alpha_{0}\in(0,1)$. ∎
This uniform estimate is the key ingredient ensuring uniform quasi-compactness of the family $(P_{\delta})_{\delta>0}$.
@@ 4. Existence and Convergence of Invariant Densities
In this section, we prove the existence of invariant densities for the perturbed operators $P_{\delta}$ and establish their convergence in $L^{1}(I)$ to the invariant density of the unperturbed system.
@@
For each $\delta>0$, the operator $P_{\delta}$ admits an invariant density $h_{\delta}\in BV_{1,1/p}(I)$ such that
@@
Let $f\in BV_{1,1/p}(I)$ with $f\geq 0$ and $\|f\|_{1}=1$. By Theorem 3.6, there exist constants $\alpha\in(0,1)$ and $C>0$, independent of $\delta$, such that
$\|P_{\delta}g\|_{1,1/p}\leq\alpha\|g\|_{1,1/p}+C\|g\|_{1}\quad\text{for all }g\in BV_{1,1/p}(I).$
Iterating and using $\|P^{k}_{\delta}f\|_{1}=\|f\|_{1}=1$ (since $f\geq 0$ and $P_{\delta}$ is Markov, so it preserves the $L^{1}(I)$ norm of nonnegative functions), we obtain for all $n\in\mathbb{N}$:
$\|P_{\delta}^{n}f\|_{1,1/p}\leq\alpha^{n}\|f\|_{1,1/p}+C\sum_{k=0}^{n-1}\alpha^{k}\|f\|_{1}\leq\|f\|_{1,1/p}+\frac{C}{1-\alpha}=:C^{\prime}.$
Define the Cesàro averages $h_{n}:=\frac{1}{n}\sum_{k=0}^{n-1}P_{\delta}^{k}f$. By convexity of the norm, $\|h_{n}\|_{1,1/p}\leq C^{\prime}$, so $\{h_{n}\}$ is bounded in $BV_{1,1/p}(I)$. Since the embedding $BV_{1,1/p}(I)\hookrightarrow L^{1}(I)$ is compact, there exist a subsequence $\{h_{n_{j}}\}$ and $h_{\delta}\in L^{1}(I)$ such that $h_{n_{j}}\to h_{\delta}$ in $L^{1}(I)$.
By lower semicontinuity of $\operatorname{var}_{1,1/p}$ with respect to $L^{1}(I)$ convergence [3],
$\operatorname{var}_{1,1/p}(h_{\delta})\leq\liminf_{j\to\infty}\operatorname{var}_{1,1/p}(h_{n_{j}})\leq C^{\prime},$
so $h_{\delta}\in BV_{1,1/p}(I)$.
@@
Taking $L^{1}(I)$ norms and using $\|P_{\delta}^{n}f\|_{1}=\|f\|_{1}=1$:
@@
Since $P_{\delta}:L^{1}(I)\to L^{1}(I)$ is a bounded linear operator, $h_{n_{j}}\to h_{\delta}$ in $L^{1}(I)$ implies $P_{\delta}h_{n_{j}}\to P_{\delta}h_{\delta}$ in $L^{1}(I)$. Combined with $\|P_{\delta}h_{n_{j}}-h_{n_{j}}\|_{1}\to 0$ and $h_{n_{j}}\to h_{\delta}$, we conclude $P_{\delta}h_{\delta}=h_{\delta}$.
Finally, since $h_{n_{j}}\to h_{\delta}$ in $L^{1}(I)$, there is a further subsequence along which $h_{n_{j}}(x)\to h_{\delta}(x)$ a.e. Since each $h_{n_{j}}\geq 0$, we get $h_{\delta}\geq 0$ a.e. Moreover, $\|h_{\delta}\|_{1}=\lim_{j\to\infty}\|h_{n_{j}}\|_{1}=1$. ∎
@@
$\sup_{g\in K}\|Q_{\delta}g-g\|_{1}\to 0\qquad\text{as }\delta\to 0.$
@@
Since $K$ is relatively compact in $L^{1}(I)$, it is totally bounded. Hence, for any $\varepsilon>0$, there exist $g_{1},\ldots,g_{N}\in L^{1}(I)$ such that for every $g\in K$ there exists $i$ with
@@
It remains to show that for each fixed $i$,
$\|Q_{\delta}g_{i}-g_{i}\|_{1}\to 0\qquad\text{as }\delta\to 0.$
Let $g_{i}^{\mathbb{T}}$ denote the periodic extension of $g_{i}$ to $\mathbb{T}$. Since $Q_{\delta}$ coincides with convolution on $\mathbb{T}$ (via periodic extension), we have
$(Q_{\delta}g_{i})(x)=(Q_{\delta}^{\mathbb{T}}g_{i}^{\mathbb{T}})(x),\qquad x\in I,$
where $Q_{\delta}^{\mathbb{T}}$ is the torus convolution with kernel $q_{\delta}(x,y)=\frac{1}{\delta}q\!\left(\frac{x-y}{\delta}\right)$ on $\mathbb{T}$. Because $q_{\delta}$ is the periodized convolution kernel with total mass $1$ and support shrinking to a point as $\delta\to 0$, it forms an approximate identity on $\mathbb{T}$; it follows that
$\|Q_{\delta}^{\mathbb{T}}g_{i}^{\mathbb{T}}-g_{i}^{\mathbb{T}}\|_{L^{1}(\mathbb{T})}\to 0.$
Restricting to $I$, we obtain
$\|Q_{\delta}g_{i}-g_{i}\|_{1}\leq\|Q_{\delta}^{\mathbb{T}}g_{i}^{\mathbb{T}}-g_{i}^{\mathbb{T}}\|_{L^{1}(\mathbb{T})}\to 0.$
@@
Assume $\tau\in\mathcal{T}([0,1];s,\varepsilon)$ is topologically mixing, so that $P_{\tau}$ admits a unique ACIM with density $h\in BV_{1,1/p}(I)$. Let $h_{\delta}$ be the invariant density of $P_{\delta}$ given by Theorem 4.1. Then $\|h_{\delta}-h\|_{1}\to 0$ as $\delta\to 0$.
@@
We define an equivalent norm on $BV_{1,1/p}(I)$ by
$\|f\|_{\mathcal{B}}:=\operatorname{var}_{1,1/p}(f)+A\|f\|_{1},\qquad A>0\text{ chosen below},$
as the strong norm, and retain $\|\cdot\|_{1}$ as the weak norm.
@@
$\operatorname{var}_{1,1/p}(P_{\delta}f)\leq\alpha\operatorname{var}_{1,1/p}(f)+C\|f\|_{1},\qquad\|P_{\delta}f\|_{1}\leq\|f\|_{1}.$
Choose $A>C/(1-\alpha)$. Adding $A$ times the second inequality to the first:
$\|P_{\delta}f\|_{\mathcal{B}}\leq\alpha\|f\|_{\mathcal{B}}+B\|f\|_{1},\qquad B:=C+A(1-\alpha),$
uniformly in $\delta>0$. Since $\|f\|_{1}\leq A^{-1}\|f\|_{\mathcal{B}}$:
$\|P_{\delta}f\|_{\mathcal{B}}\leq M\|f\|_{\mathcal{B}},\quad M:=\alpha+\frac{B}{A}=\alpha+\frac{C}{A}+(1-\alpha)=1+\frac{C}{A}<\infty,$
@@
Step 2: Perturbation smallness $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}(I)}\to 0$. Write $P_{\delta}f-P_{\tau}f=(Q_{\delta}-I)P_{\tau}f$. Since $P_{\tau}$ satisfies the same Lasota–Yorke inequality [2], it maps the unit ball of $(BV_{1,1/p}(I),\|\cdot\|_{\mathcal{B}})$ into a set bounded in $BV_{1,1/p}(I)$. By compactness of $BV_{1,1/p}(I)\hookrightarrow L^{1}(I)$, the set
@@
and hence $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}(I)}\to 0$.
@@
Quasi-compactness. By the Lasota–Yorke inequality for $P_{\tau}$ and the compact embedding $BV_{1,1/p}(I)\hookrightarrow L^{1}(I)$, the Ionescu–Tulcea–Marinescu theorem [5] implies that $P_{\tau}$ is quasi-compact on $(BV_{1,1/p}(I),\|\cdot\|_{\mathcal{B}})$, with essential spectral radius $r_{\mathrm{ess}}(P_{\tau})\leq\alpha<1$.
Simplicity of eigenvalue $1$. Let $g\in BV_{1,1/p}(I)$ with $P_{\tau}g=g$. Write $g=g^{+}-g^{-}$ with $g^{\pm}\geq 0$. By positivity of $P_{\tau}$: $P_{\tau}|g|\geq|P_{\tau}g|=|g|$. Since $P_{\tau}$ is a Markov operator, $\int P_{\tau}|g|\,dm=\int|g|\,dm$. Since $P_{\tau}|g|\geq|g|$ a.e. and both functions have the same integral, $P_{\tau}|g|=|g|$ a.e. Adding this to $P_{\tau}g=g$:
@@
so $P_{\tau}g^{+}=g^{+}$ a.e., and likewise $P_{\tau}g^{-}=g^{-}$ a.e. If $g^{+}\not\equiv 0$, then $g^{+}/\|g^{+}\|_{1}$ is a normalized nonnegative fixed point of $P_{\tau}$; by uniqueness of the ACIM (standard under topological mixing [6, 8]; see Remark 1 below), $g^{+}=c^{+}h$. Similarly $g^{-}=c^{-}h$. Hence $g=(c^{+}-c^{-})h$, and $\ker(P_{\tau}-I)$ is one-dimensional.
Isolation. Since $r_{\mathrm{ess}}(P_{\tau})\leq\alpha<1$, all eigenvalues of $P_{\tau}$ on the unit circle are isolated with finite multiplicity. The eigenvalue $1$ is simple by the above, hence isolated. Its spectral projection is $\Pi f=\bigl(\int f\,dm\bigr)h$, which has rank one.
Step 4: Application of Keller–Liverani. Steps 1–3 verify the hypotheses of [4, Theorem 1]:
@@
    Uniform Lasota–Yorke: $\|P_{\delta}f\|_{\mathcal{B}}\leq\alpha\|f\|_{\mathcal{B}}+B\|f\|_{1}$, with $\alpha\in(0,1)$ and $B<\infty$ independent of $\delta$.
@@
    Perturbation smallness: $\|P_{\delta}-P_{\tau}\|_{\mathcal{B}\to L^{1}(I)}\to 0$.
@@
    Spectral isolation: eigenvalue $1$ of $P_{\tau}$ is isolated and simple, with rank-one spectral projection $\Pi$.